# Valid F-screening in linear regression

### v1

@@ B.2 Confidence intervals
unconditionally follows an $F_{1,n-p-1}$ distribution under $H_{0}^{j}(b)$ (Seber & Lee 2003, §4.3). Thus the analogous selective p-value to that in (10) is
$(\displaystyle{p_{H_{0}^{j}(b)\mid E}}(y):=Pr_{\beta_{j}=b}\Bigg{(}$$\frac{(Y-bX_{j})^{\top}(P_{X}-P_{X_{-j}})(Y-bX_{j})}{Y^{\top}(I_{n-1}-P_{X})Y}\geq r^{\prime}(y)\Big{|}Y\in E(y)\Bigg{)},$ (B.10)
@@
$(\displaystyle{p_{H_{0}^{j}(b)\mid E}}(y)=Pr_{\beta_{j}=b}\Bigg{(}$$\frac{(Y-bX_{j})^{\top}(P_{X}-P_{X_{-j}})(Y-bX_{j})}{Y^{\top}(I_{n-1}-P_{X})Y}\geq r^{\prime}(y)$
@@
where $Z\sim\chi^{2}_{n-p-1}$ and $G\sim\mathcal{N}(U_{j}^{\top}X_{j}b,\sigma^{2}).$ The advantage of (3) over (B.2) is that the former can be computed using only univariate variables, whereas the latter involves the generation of $(n-1)$-dimensional random vectors in each Monte Carlo replicate.
@@ E.1 A useful lemma
[…]
@@
To prove claims 1-3, we recall that if $Y\sim\mathcal{N}_{n-1}(\mu,\sigma^{2}I_{n-1})$ and $P$ a projection matrix of rank $r$, then $Y^{\top}PY/\sigma^{2}\sim\chi^{2}_{r}(\lambda)$ where $\lambda=\mu^{\top}P\mu/(2\sigma^{2})$ is the noncentrality parameter. Using this result, to prove part 1 of Lemma E.1, observe that $\frac{1}{\sigma}Y\sim\mathcal{N}_{n-1}\left(\frac{1}{\sigma}X\beta,I_{n-1}\right)$ and $I_{n-1}-P_{X}$ is a projection matrix of rank $n-p-1$. It then follows that
@@
$\frac{1}{\sigma^{2}}Y^{\top}(P_{X}-P_{X_{-M}})Y\sim\chi^{2}_{m}\left(\frac{(X_{M}\beta_{M})^{\top}(I_{n-1}-P_{X_{-M}})X_{M}\beta_{M}}{2\sigma^{2}}\right).$
@@
Finally, to prove the third statement, note that $P_{X_{-M}}$ is a projection matrix of rank $p-m$. Hence,
$\frac{1}{\sigma^{2}}Y^{\top}P_{X_{-M}}Y\sim\chi^{2}_{p-m}\left(\frac{1}{2\sigma^{2}}(X\beta)^{\top}P_{X_{-M}}X\beta\right).$
@@
With $E_{1}$ from (6) and $E_{2}(y)$ from (8), from the law of total expectation we obtain that
@@
From part 4 of Lemma E.1, we have that $Y^{\top}(P_{X}-P_{{X}_{-M}})Y$, $Y^{\top}(I_{n-1}-P_{X})Y$, and $Y^{\top}P_{{X}_{-M}}Y$ are mutually independent. Importantly, this implies that $Y^{\top}P_{{X}_{-M}}Y$ is independent of any function of $Y^{\top}(P_{X}-P_{{X}_{-M}})Y$ and $Y^{\top}(I_{n-1}-P_{X})Y$. Therefore $Y^{\top}P_{{X}_{-M}}Y$ is independent of both $Y^{\top}(P_{X}-P_{X_{-M}})Y/Y^{\top}(I_{n-1}-P_{X})Y$ and $Y^{\top}(P_{X}-P_{X_{-M}})Y-cY^{\top}(I_{n-1}-P_{X})Y$, as well as their intersection. For $d$ be defined as in (13), it follows that
@@
By Lemma E.1 parts 1 and 2, $\tfrac{Y^{\top}(P_{X}-P_{{X}_{-M}})Y}{Y^{\top}(I_{n-1}-P_{X})Y}$ is distributed as $W/Z$ and $Y^{\top}(P_{X}-P_{{X}_{-M}})Y-cY^{\top}(I_{n-1}-P_{X})Y$ is distributed as $\sigma^{2}W-c\sigma^{2}Z$ where $W\sim\chi_{m}^{2}$ and $Z\sim\chi_{n-p-1}^{2}$. Furthermore, by part 4 of Lemma E.1, $Y^{\top}(P_{X}-P_{{X}_{-M}})Y$ and $Y^{\top}(I_{n-1}-P_{X})Y$ are independent. Using the representation of ${p_{H_{0}^{M}\mid E}}$ introduced in equation (E.1), the result follows. ∎

### v2

@@ B.2 Confidence intervals
unconditionally follows an $F_{1,n-p-1}$ distribution under $H_{0}^{j}(b)$ (Seber & Lee, 2003, §4.3). Thus the analogous selective p-value to that in (9) is
${p_{H_{0}^{j}(b)\mid E}}(y):=Pr_{\beta_{j}=b}\Bigg($ $\frac{(Y-bX_{j})^{\top}(P_{X}-P_{X_{-j}})(Y-bX_{j})}{Y^{\top}(I_{n-1}-P_{X})Y}\geq r^{\prime}(y)\Big|Y\in E(y)\Bigg),$ (B.10)
@@
${p_{H_{0}^{j}(b)\mid E}}(y)=Pr_{\beta_{j}=b}\Bigg($ $\frac{(Y-bX_{j})^{\top}(P_{X}-P_{X_{-j}})(Y-bX_{j})}{Y^{\top}(I_{n-1}-P_{X})Y}\geq r^{\prime}(y)$
@@
where $Z\sim\chi^{2}_{n-p-1}$ and $G\sim\mathcal{N}(U_{j}^{\top}X_{j}b,\sigma^{2}).$ The advantage of (4) over (B.2) is that the former can be computed using only univariate variables, whereas the latter involves the generation of (n1)(n-1)-dimensional random vectors in each Monte Carlo replicate.
@@ E.1 A useful lemma
$Y^{\top}(I_{n-1}-P_{X})Y=\left\lVert(I_{n-1}-P_{X})Y\right\rVert_{2}^{2}\hskip 8.53581pt\text{and}\hskip 8.53581ptY^{\top}P_{X_{-M}}Y=\left\lVert P_{X_{-M}}Y\right\rVert_{2}^{2}.$
@@
To prove claims 1-3, we recall that if $Y\sim\mathcal{N}_{n-1}(\mu,\sigma^{2}I_{n-1})$ and $P$ a projection matrix of rank $r$, then $Y^{\top}PY/\sigma^{2}\sim\chi^{2}_{r}(\lambda)$ where $\lambda=\mu^{\top}P\mu/(2\sigma^{2})$ is the noncentrality parameter. Using this result, to prove part 1 of Lemma E.1, observe that $\frac{1}{\sigma}Y\sim\mathcal{N}_{n-1}\left(\frac{1}{\sigma}X\beta,I_{n-1}\right)$ and $I_{n-1}-P_{X}$ is a projection matrix of rank np1n-p-1. It then follows that
@@
$\frac{1}{\sigma^{2}}Y^{\top}(P_{X}-P_{X_{-M}})Y\sim\chi^{2}_{m}\left(\frac{(X_{M}\beta_{M})^{\top}(I_{n-1}-P_{X_{-M}})X_{M}\beta_{M}}{2\sigma^{2}}\right).$ (E.1)
@@
Finally, to prove the third statement, note that $P_{X_{-M}}$ is a projection matrix of rank pmp-m. Hence,
$\frac{1}{\sigma^{2}}Y^{\top}P_{X_{-M}}Y\sim\chi^{2}_{p-m}\left(\frac{1}{2\sigma^{2}}(X\beta)^{\top}P_{X_{-M}}X\beta\right).$ (E.2)
@@
Because $Y\in E\implies Y\in E_{1}$ for $E_{1}$ defined in (6) and $E(y)$ defined in (8), when we apply the law of total expectation, we obtain
@@
From part 4 of Lemma E.1, we have that $Y^{\top}(P_{X}-P_{{X}_{-M}})Y$, $Y^{\top}(I_{n-1}-P_{X})Y$, and $Y^{\top}P_{{X}_{-M}}Y$ are mutually independent. Importantly, this implies that $Y^{\top}P_{{X}_{-M}}Y$ is independent of any function of $Y^{\top}(P_{X}-P_{{X}_{-M}})Y$ and $Y^{\top}(I_{n-1}-P_{X})Y$. Therefore $Y^{\top}P_{{X}_{-M}}Y$ is independent of both $Y^{\top}(P_{X}-P_{X_{-M}})Y/Y^{\top}(I_{n-1}-P_{X})Y$ and $Y^{\top}(P_{X}-P_{X_{-M}})Y-cY^{\top}(I_{n-1}-P_{X})Y$, as well as their intersection. For $d$ defined as in (11), it follows that
@@
By Lemma E.1 parts 1 and 2, under $H_{0}^{1:p}$, $\tfrac{Y^{\top}(P_{X}-P_{{X}_{-M}})Y}{Y^{\top}(I_{n-1}-P_{X})Y}$ is distributed as $W/Z$ and $Y^{\top}(P_{X}-P_{{X}_{-M}})Y-cY^{\top}(I_{n-1}-P_{X})Y$ is distributed as $\sigma^{2}W-c\sigma^{2}Z$ where $W\sim\chi_{m}^{2}$ and $Z\sim\chi_{n-p-1}^{2}$. Furthermore, by part 4 of Lemma E.1, $Y^{\top}(P_{X}-P_{{X}_{-M}})Y$ and $Y^{\top}(I_{n-1}-P_{X})Y$ are independent. Using the representation of $p_{H_{0}^{M}\mid E}$} introduced in (E.3), the result follows. ∎
@@
We begin with two lemmas.
Lemma E.3.
Consider a fixed sequence of design matrices $X_{n}\in\mathbb{R}^{n\times p}$ and corresponding realizations of the linear model $Y_{n}=X_{n}\beta+\epsilon_{n}$ from (1). Suppose that $\frac{1}{n}X_{n}^{\top}X_{n}\to Q$ for some positive definite $Q$. Then $\exists\delta>0$ and $N>0$ such that $\Pr(W-c_{n}Z_{n}\geq d(Y_{n})/\sigma^{2})\geq\delta$ for all $n>N$, where $W\sim\chi^{2}_{m}$, $Z_{n}\sim\chi^{2}_{n-p-1}$, $c_{n}=\frac{p}{n-p-1}F^{-1}_{p,n-p-1}\left(1-\alpha_{0}\right)$, and $d(Y_{n})=-Y_{n}^{\top}P_{\left(X_{n}\right)_{-M}}Y_{n}$, with $W$, $Z_{n}$, and $d(Y_{n})$ independent.
Proof of Lemma E.3.
In what follows, we suppress subscripts on $X$ involving $n$ to enhance readability.
To show that $\Pr(W-c_{n}Z_{n}\geq d(Y_{n})/\sigma^{2})>0$ as $n\to\infty$, first observe that $c_{n}\cdot Z_{n}$ converges in probability to $(\chi^{2}_{p})^{-1}(1-\alpha_{0})$ as $n\to\infty$. On the other hand, $d(Y_{n})$ is a scaled chi-square random variable with pmp-m degrees of freedom and non-centrality parameter $\tfrac{1}{2\sigma^{2}}(X\beta)^{\top}P_{X_{-M}}(X\beta)$. We treat the cases where $\beta\neq 0$ and $\beta=0$ separately.
Case 1: Suppose that $\beta\neq 0$. Then the non-centrality parameter of $d(Y_{n})$ is possibly non-zero. By our assumption that $\frac{1}{n}X^{\top}X\to Q$ deterministically for some positive definite matrix $Q\in\mathbb{R}^{p\times p}$, it follows that $\frac{1}{n}X^{\top}X_{-M}\to Q_{\cdot,-M}$ and $\frac{1}{n}X_{-M}^{\top}X_{-M}\to Q_{-M,-M}$ for $Q_{\cdot,-M}\in\mathbb{R}^{p\times(p-m)}$ and positive definite $Q_{-M,-M}\in\mathbb{R}^{(p-m)\times(p-m)}$. Hence,
$\tfrac{1}{n}\beta^{\top}X^{\top}P_{X_{-M}}X\beta$ $=\beta^{\top}\left(\tfrac{1}{n}X^{\top}X_{-M}\right)\left(\tfrac{1}{n}X_{-M}^{\top}X_{-M}\right)^{-1}\left(\tfrac{1}{n}X_{-M}^{\top}X\right)\beta$
$\to\beta^{\top}Q_{\cdot,-M}(Q_{-M,-M})^{-1}Q_{\cdot,-M}^{\top}\beta.$
Thus, when $\beta\neq 0$, $d(Y_{n})$ is distributed as a scaled chi-square random variable with non-centrality parameter that is $O(n)$. The sequence of expectations of $d(Y_{n})/n$ converges and its variance is $O(1/n).$ Thus by Chebyshev’s inequality, it follows that $\phi_{n}:=d(Y_{n})/n\overset{P}{\to}\phi^{*}$ for some fixed, negative $\phi^{*}$, where $\overset{P}{\to}$ denotes convergence in probability.
Therefore, for fixed $\varepsilon>0$ and $\eta>0$, we have that $\Pr(|\phi_{n}-\phi^{*}|\geq\varepsilon)<\eta$ and $\Pr(|c_{n}Z_{n}-(\chi^{2}_{p})^{-1}(1-\alpha)|\geq\varepsilon)<\eta$ for sufficiently large $n$. For such an $n$, define
$G_{n}(\varepsilon)=\{(\phi_{n},Z_{n}):|\phi_{n}-\phi^{*}|<\varepsilon,\;|c_{n}Z_{n}-(\chi^{2}_{p})^{-1}(1-\alpha)|<\varepsilon\},$
so that $\Pr(G_{n}^{c}(\varepsilon))\leq 2\eta$. Hence
$\Pr\bigl(W-c_{n}Z_{n}\geq d(Y_{n})/\sigma^{2}\bigr)$ $=\Pr\bigl(W-c_{n}Z_{n}\geq n\phi_{n}/\sigma^{2}\bigr)$
$\geq\Pr\bigl(W\geq c_{n}Z_{n}+n\phi_{n}/\sigma^{2},\;G_{n}(\varepsilon)\bigr)$
$\geq\Pr\bigl(W\geq(\chi^{2}_{p})^{-1}(1-\alpha)+\varepsilon+n(\phi^{*}+\varepsilon)/\sigma^{2},\;G_{n}(\varepsilon)\bigr)$
$\geq\Pr\bigl(W\geq(\chi^{2}_{p})^{-1}(1-\alpha)+\varepsilon+n(\phi^{*}+\varepsilon)/\sigma^{2}\bigr)-\Pr\bigl(G_{n}^{c}(\varepsilon)\bigr)$
$\geq\Pr\bigl(W\geq(\chi^{2}_{p})^{-1}(1-\alpha)+\varepsilon+n(\phi^{*}+\varepsilon)/\sigma^{2}\bigr)-2\eta.$
Note that $W\sim\chi^{2}_{m}$ has support $0,\infty)$ and, recalling that $\phi^{*}$ is negative, $(\chi^{2}_{p})^{-1}(1-\alpha)+\varepsilon+n(\phi_{n}^{*}+\varepsilon)/\sigma^{2}$ approaches -$\infty$ as $n\to\infty$. Since $\eta$ can be arbitrarily close to 0, we conclude that $\Pr\left(W-c_{n}Z_{n}\geq d(Y_{n})/\sigma^{2}\right)\to 1$ as $n\to\infty$, and hence is bounded away from 0 in the limit.
Case 2: Under the global null $H_{0}^{1:p}:\beta=0,$ the non-centrality parameter of $d(Y_{n})$ is 0. In this case, the distribution of $d(Y_{n})$ does not change with $n$. For fixed $\eta>0$, $\varepsilon>0$, we have that $\Pr(|c_{n}Z_{n}-(\chi^{2}_{p})^{-1}(1-\alpha_{0})|\geq\varepsilon)<\eta$ for sufficiently large $n$, and we can choose $B_{\eta}$ such that $\Pr(|d(Y_{n})|\geq B_{\eta})<\eta$. Then define
$G_{n}(\varepsilon,B_{\eta})=\{(d(Y_{n}),Z_{n}):|d(Y_{n})|<B_{\eta},\;|c_{n}Z_{n}-(\chi^{2}_{p})^{-1}(1-\alpha_{0})|<\varepsilon\},$
so that $\Pr(G_{n}^{c}(\varepsilon,B_{\eta}))\leq 2\eta$. Hence, similar to case 1,
$\Pr\bigl(W-c_{n}Z_{n}\geq d(Y_{n})/\sigma^{2}\bigr)\geq\Pr\bigl(W\geq(\chi^{2}_{p})^{-1}(1-\alpha_{0})+\varepsilon+B_{\eta}/\sigma^{2}\bigr)-2\eta.$
Letting $\eta$ approach 0, we conclude that $\Pr\left(W-c_{n}Z_{n}\geq d(Y_{n})/\sigma^{2}\right)>0$ in the limit since $(\chi^{2}_{p})^{-1}(1-\alpha_{0})+\varepsilon+B_{\eta}/\sigma^{2}$ is a fixed finite quantity that does not change (and in particular does not grow) with $n$.
Lemma E.4.
Suppose $f:\mathbb{R}^{n}\to\mathbb{R}$ is a function that satisfies $\Pr(f(Y_{n})>\varepsilon)\to 0$ for every $\varepsilon>0$. Then $\Pr(f(Y_{n})>\varepsilon\mid Y_{n}\in E_{1})\to 0$ for every $\varepsilon>0$.
Proof of Lemma E.4.
Note that $\Pr_{Y_{n}}(Y_{n}\in E_{1})\geq\alpha_{0}$ where $\alpha_{0}>0$ is the level of the F-test of $H_{0}^{1:p}$ in Step 1 of Box 1. Thus
$\Pr_{Y_{n}}\left(f(Y_{n})>\varepsilon\right)\to 0$
$\quad\implies\Pr_{Y_{n}}\left(f(Y_{n})>\varepsilon\mid Y_{n}\in E_{1}\right)\Pr_{Y_{n}}\left(Y_{n}\in E_{1}\right)$
$\quad\quad\quad\quad\quad+\Pr_{Y_{n}}\left(f(Y_{n})>\varepsilon\mid Y_{n}\notin E_{1}\right)\Pr_{Y_{n}}\left(Y_{n}\notin E_{1}\right)\to 0$
$\quad\implies\Pr_{Y_{n}}\left(f(Y_{n})>\varepsilon\mid Y_{n}\in E_{1}\right)\Pr_{Y_{n}}\left(Y_{n}\in E_{1}\right)\to 0$
$\quad\implies\Pr_{Y_{n}}\left(f(Y_{n})>\varepsilon\mid Y_{n}\in E_{1}\right)\to 0.$
We now continue with the proof of Proposition 2.2.